### v2

@@ 3.1.2 Prompt Tuning
In prompt tuning [89], task-specific "soft prompts" are added to the input data. These are opaque learnable tensors concatenated with the input embedding as shown in figure 9. Throughout training, only the prompts are modified for the specified task. This strategy is extremely beneficial for multitasking and continual learning. [197] introduce Knowledge-guided Context Optimization (KgCoOp) to enhance the generalization capability of prompt tuning for VLMs. Their strategy reduces the discrepancy between learned soft prompts and hard prompts, thus reducing the likelihood of forgetting important general knowledge. By integrating this optimization constraint with contrastive learning, their method improves performance on unseen classes while also maintaining computational efficiency. [187] propose a dual-modality prompt tuning (DPT) approach that optimizes both visual and textual prompts simultaneously to adapt vision-language models to downstream tasks. They also introduce class -aware visual prompts, generated by cross-attention between textual features and image tokens, to improve task-specific feature extraction. To address the misalignment between text and image embeddings in VLMs, [22] develop DAPT to optimize feature distributions by maximizing inter-class separation and minimizing intra-class variability. This approach improves generalization in few-shot learning and domain adaptation.
3.1.3 Adapter-based Transfer Leraqning
@@
$X\in\mathbb{R}^{a\times a}\rightarrow W_{d}X\in\mathbb{R}^{a\times b}\rightarrow W_{u}(W_{d}X)\in\mathbb{R}^{a\times a}.$ (12)
@@ 3.1.4 Prefix Tuning
Prefix Tuning [97] introduces task-specific continuous vectors into each layer of the transformer architecture. Prefixes are abstract representations concatenated with input embeddings in each layer. These vectors are optimized during fine-tuning and act as soft prompts that steer the model toward task-specific behavior without updating the backbone parameters. [160] introduce a mapping network that transforms visual features into token embeddings and employs prefix tuning to to refine query  representations in each transformer layer for open-ended medical visual question answering (VQA). [159] propose a vision-language storytelling framework that integrates a trainable transformer mapping network and a multimodal contrastive objective to enhance coherence in generated narratives. Prefix tuning is employed to align CLIP-extracted visual features with a frozen GPT-2 [147] model, facilitating efficient vision-language alignment. Prefixes are introduced in personalized image captioning to map visual features with user context, enabling a frozen GPT-2 model to provide personalized captions [179].
@@ 4 Runtime Optimizations
Runtime optimization refers to techniques that dynamically adjust computations during  inference to improve computational efficiency, memory use, and energy consumption without requiring retraining. Unlike static optimizations such as pruning or quantization, runtime methods adapt model execution on-the-fly based on input complexity, hardware constraints, or task-specific requirements. VLMs are computationally intensive, necessitating significant processing for multimodal interactions. Their deployment on  resource-constrained devices is challenging due to the high memory, power, and latency demands. Runtime optimization mitigates these issues by reducing unnecessary computations and accelerating inference. The datasets evaluated remain consistent with those in previous sections, with top-k accuracy [40, 123] being the primary evaluation metric. GPU consumption is also lower when compared to other techniques previously described, as computations are adjusted dynamically during inference.
@@ 4.2 Test-Time Adaptation (TTA)
$p(y\mid x)=\arg\max_{y}\mathbb{E}_{q(\theta)}[p(y\mid x;\theta)].$ (13)
@@ 4.2.2 Test-Time Prompt Tuning (TPT)
We compare the various paradigms and techniques for engineering efficient VLMs in Table 2. The summary provides an overview of the computational efficiency and memory requirements of the methods described in sections two through four and the best use cases for each technique. Pre-deployment techniques give the best efficiency in terms of inference time and storage because they shrink the model before deployment, leading to a smaller model on the disk. Efficient fine-tuning techniques do not speed up inference nor reduce storage because they leave the entire model intact and only add adapters or prompts. Runtime optimizations eliminate redundant inference computations to deliver some speedup while preserving the majority of model parameters.
Table 2: Comparison of computational efficiency and memory requirement across various techniques. $\bullet$ indicates best performance; $\bullet$ indicates medium performance; $\bullet$ indicates worst performance.
Paradigm Technique Compute Efficiency Memory Requirements Use Case
Post-Training Quantization $\bullet$ $\bullet$ Reduce model size and improve latency at deployment. Useful to deploy on low-memory devices at the cost of accuracy.
Pre-Deployment Structured Pruning $\bullet$ $\bullet$ Remove redundant weights for sparse model deployment. Useful if CPU/GPU supports sparse matrix ops.
Knowledge Distillation $\bullet$ $\bullet$ Train a smaller student for fast, lightweight inference. Useful for application specific deployment
Low-Rank Adaptation (LoRA) $\bullet$ $\bullet$ Low-resource fine-tuning without changing base weights. Requires server to fine-tune the model
Efficient Fine-Tuning Prompt Tuning $\bullet$ $\bullet$ Efficient task adaptation with minimal updates. Requires server to fine-tune the model unless on tiny VLMs
Adapter based Transfer Learning $\bullet$ $\bullet$ Reusable modules across tasks without retraining full model.
Token Reduction $\bullet$ $\bullet$ Speed up inference by processing fewer tokens.
Runtime Optimization Test-Time Adaptation $\bullet$ $\bullet$ Dynamically adapt to domain shifts or out-of-distribution data.
Test-Time Prompt Tuning $\bullet$ $\bullet$ Apply small prompts at runtime for improved generalization.
@@ 5 Privacy-preserving Distributed VLM
Privacy-preserving distributed VLM ensures that model training occurs across multiple devices or servers without sharing raw data. The key benefits include enhanced data security and a decreased risk of data leakage. The datasets and evaluation metrics used remain the same as in previous sections. However, computational resources differ depending on the client devices. For instance, a Jetson Nano has 4GB RAM, while a Jetson Orin offers 8GB RAM, reflecting one example of such variation. Below, we discuss key techniques such as Federated Learning (FL) 5.1 and Split Learning 5.2.
@@ 6.5 MiniVLM
DeepSpeed  [^1] https://www.deepspeed.ai/ Zero and sparse training optimizations Distributed fine-tuning and model parallelism Large VLMs (e.g., LLaVA variants, Flamingo, MiniGPT-4, etc.) Not directly (requires custom FL loop on top of DeepSpeed modules)
@@
6.6 LightCLIP
6.7 NanoVLMs
6.8 Moondream
6.9 Xmodel-VLM
Table 4: Efficient models with parameter size, methods used, and advantages.
Model/Method Name Parameters Method(s) Advantage
CLIP-LoRA [207] 149M total / $\sim$16M trainable LoRA on query, key, value matrices of both vision and text encoders (rank=2). Outperforms or matches full fine-tuning performance across 11 few-shot classification datasets.
PromptKD [100] $\approx$ 340M (student) Prompt-based distillation using pre-stored class vectors and KL alignment of image logits with teacher CLIP. Improves zero-shot accuracy by 12% (HM) over CLIP across 11 datasets.
KD-VLP [113] $\approx$ 138M (student) Object-guided masked vision modeling & phrase-region alignment using knowledge distillation from external object detectors. Achieves comparable retrieval and reasoning performance with significantly fewer parameters.
LoRA-Sparse [158] $\approx$ 600M (upper bound estimate of trainable parameters) Applies low-rank linear projection layers for sparse attention approximation. Maintains baseline accuracy with 70% fewer trainable parameters.
MedViLL [137] $\approx$ 110M Joint masked language modeling & image-report alignment using BERT with multimodal attention masking. Outperforms baselines on both VLU and VLG tasks with strong generalization to unseen data.
TDA [77] $\approx$ 102M (RN50) / $\approx$ 149M (ViT-B/16); doesn’t add trainable weights beyond CLIP Training-free dynamic adapter using dual key-value caches for efficient test-time adaptation of CLIP. Achieves +1.54% accuracy gain over baseline CLIP while reducing test time to just 16 minutes without backpropagation.
FastVLM [172] 125M (vision) + LLM (e.g. 0.5B / 1.2B / 7B) Combines a hybrid FastViTHD encoder with reduced visual tokens for efficient high-resolution vision-language modeling. Achieves 3.2× improvement in time-to-first-token (TTFT) while maintaining similar performance on VLM benchmarks.
VisualBERT [94] 110M Jointly embeds text and object-level image features using a unified Transformer with visually-grounded pretraining on caption data. Outperforms Pythia baselines on VQA with 70.80% (Test-Dev) and 71.00% (Test-Std) accuracy using a simpler architecture.
Furthermore, Table 3 outlines various frameworks and libraries. This table provides readers and researchers with a comprehensive overview of each library/framework’s capabilities to help them choose the right tool. Table 4 summarizes the parameter counts, methodologies, and advantages of various compact VLMs. Although larger models often correlate with better performance, the primary intent of this table is to highlight that models with significantly fewer parameters (often in the low million range) can still outperform baseline models while offering benefits in speed and memory efficiency. This performance gain stems largely from selecting the most appropriate method for the task. For example, rows 1–3 demonstrate that techniques like LoRA and knowledge distillation enable competitive or superior performance with substantial reductions in parameter size. However, these techniques depend heavily on the objectives: For instance, Knowledge distillation yields compact student models but sacrifices the teacher’s representational richness, rendering them more suitable for specialized tasks than for multitask scenarios. Fine-tuning techniques such as LoRA are chosen for efficient adaptation with minimal additional parameters, while pre-deployment techniques such as knowledge distillation is used to transfer a large model’s full predictive behavior into a compact student network. This table demonstrates how various technqiues decsribed above yield compact VLMs that match or exceed baseline performance with far fewer parameters, while detailing their specific methods.
@@ 7 Analysis and Insights
In this section, we 1 analyze the impact of efficiency-driven techniques on accuracy, latency, and memory usage, and 2 present the performance of several VLMs across various multimodal benchmarks.
7.1 Impact of Efficiency-Boosting Techniques on Accuracy, Memory, and Latency
Our evaluation is based on two models: 1 blip-vqa-base [93] ($\sim 385M$ parameters) and 2 vilt-b32-finetuned-vqa [84] ($\sim 87M$ parameters). All results are reported using an NVIDIA RTX A6000 GPU.We use the Balanced Binary Abstract Scenes [216] dataset which is a visual question answering (VQA) dataset, where each sample consists of an abstract scene image, a natural language question, and multiple human-annotated answers. Evaluation is based on the official VQA accuracy metric, which grants full credit if at least three out of ten annotators provide the same answer as the model. We also measure average inference latency and memory consumption to assess runtime efficiency. Notably, the observed trends remain consistent across all models, provided the parameter count falls within the range of these two models. This section provides insights specifically for pre-deployment techniques such as quantization, pruning, and low-rank approximation. We focus on these methods because they are fundamental for optimizing models before deployment, directly impacting model size and performance. Although efficient fine-tuning, runtime optimizations, and privacy-preserving methods are important, they are often model- or task-specific and involve different challenges and considerations that extend beyond the scope of this survey. Our analysis provides insights into the accuracy vs. efficiency trade-off, and emphasizes the importance of selecting the right technique.
@@ 7.1 Impact of Efficiency-Boosting Techniques on Accuracy, Memory, and Latency
We conducted FP16 inference on GPU with torch.amp.autocast and INT8 inference using dynamic quantization of linear layers. Quantization (figure 14) has a notable impact on accuracy, latency, and memory consumption on both blip-vqa-base and vilt-b32-finetned-vqa. For blip-vqa-base, transitioning from FP32 to INT8 results in a 0.85% accuracy drop for 50 samples whereas vilt-b32-finetned-vqa experiences an 8.07% reduction indicating that ViLT is more sensitive to INT8 quantization. We also increase the number of samples to 125 to observe the effect on accuracy and memory consumption. Increasing the sample size from 50 to 125 causes a 1.36% accuracy drop in FP32 and a 1.54% drop in INT8 for blip-vqa-base, while vilt-b32-finetned-vqa sees a 7.1% decrease in FP32 and a 0.09% reduction in INT8. However, the accuracy drop for vilt-b32-finetuned-vqa from FP32 to INT8 is steeper (4.6% on average) compared to blip-vqa-base (0.94%), suggesting that vilt-b32-finetuned-vqa benefits more from higher precision, but suffers greater degradation from aggressive quantization. In terms of memory consumption, quantization significantly reduces memory requirements across both models. For blip-vqa-base, moving from FP32 to INT8 results in a 53.8% reduction (747.8MB to 345.2MB) for 50 samples and a 43.1% reduction (800MB to 455MB) for 125 samples, while vilt-b32-finetuned-vqa achieves a 57.2% reduction (586.3MB to 251.1MB) for 50 samples and 44.2% (664.4MB to 370.6MB) for 125 samples, demonstrating that INT8 provides substantial memory savings, with vilt-b32-finetuned-vqa benefiting slightly more. Across all quantization levels, blip-vqa-base consistently consumes more memory than vilt-b32-finetuned-vqa, with a difference of 161.5MB in FP32 (50 samples) and 94.1MB in INT8 (50 samples). The blip-vqa-base maintains a more stable accuracy across sample sizes but consistently requires more memory. The latency per sample remains 0.0877s, 0.0757s, and 0.0701s across FP32, FP16, and INT8 for blip-vqa-base and 0.0330s, 0.0321s, and 0.0311s for vilt-b32-finetuned-vqa, with INT8 being the fastest in both cases. For FP32, we run the models with full 32-bit precision. For FP16, mixed precision is enabled using autocasting to boost speed while maintaining accuracy. For INT8, a copy is dynamically quantized on its linear layers to shrink the model and speed up inference.
@@
We applied global unstructured L1 pruning to all linear layers at 20%, 40%, and 60% sparsity to evaluate the impact of weight sparsification. Pruning (figure 15) significantly impacts accuracy and memory consumption across both models. For blip-vqa-base, accuracy drops by 0% at 20% pruning, 4.24% at 40% pruning and 11.9% at 60% pruning for 50 samples. The vilt-b32-finetuned-vqa model follows a similar trend, but at 60% pruning, accuracy drops to 0%, indicating extreme sensitivity to high sparsity levels. In terms of memory savings, blip-vqa-base reduces by 16.3% at 20% pruning, 22.3% at 40% pruning, and 27% at 60% pruning (50 samples), while vilt-b32-finetuned-vqa achieves reductions of 1.7%, 15.4%, and 18.5%, respectively. For 125 samples, blip-vqa-base saves 12.3%, 20.7%, and 24.9%, whereas vilt-b32-finetuned-vqa reduces by 4.2%, 13.4%, and 21%. The results show that vilt-b32-finetuned-vqa maintains better accuracy at lower pruning levels but is highly sensitive to aggressive pruning. The latency per sample for blip-vqa-base at 20%, 40%, and 60% pruning is 0.0569s, 0.0561s, and 0.0559s, respectively, while for vilt-b32-finetuned-vqa, it is 0.0148s, 0.0141s, and 0.0139s at the same pruning levels. We observe that pruning results in faster inference per sample compared to quantization. We implement pruning by scanning all model modules to find linear layers and then apply global unstructured L1 pruning with a set sparsity ratio. After pruning, the extra reparameterization is removed, making the pruned model ready for inference.
@@
We applied low-rank approximation to every linear layer by truncating singular values at rank ratios of 0.8, 0.6, and 0.4. Low-rank approximation (figure 16) significantly affects accuracy and memory consumption across both models. Decreasing the rank from k = 0.8 to k = 0.4 increases compression, leading to lower memory consumption and latency. However, this also results in a drop in accuracy. For blip-vqa-base, accuracy decreases by 15.8% at k = 0.8, 27.8% at k = 0.6, and drops to 0% at k = 0.4 for 50 samples. The vilt-b32-finetuned-vqa exhibits a more gradual decline, with accuracy dropping by 4% at k = 0.8, 6.2% at k = 0.6, and 27.1% at k = 0.4, indicating that vilt-b32-finetuned-vqa is more resilient to LRA compared to blip-vqa-base. In terms of memory savings, blip-vqa-base reduces memory consumption by 140 MB at k = 0.8, 160.7 MB at k = 0.6, and 208.6 MB at k = 0.4 (50 samples), while vilt-b32-finetuned-vqa achieves reductions of 7.2 MB, 63.2 MB, and 86.4 MB, respectively. For 125 samples, blip-vqa-base saves 65.4 MB, 100.2 MB, and 199.3 MB, whereas vilt-b32-finetuned-vqa reduces by 64 MB, 83.8 MB, and 150 MB. This indicates that while low-rank approximation provides substantial memory savings, blip-vqa-base is highly sensitive to aggressive rank reduction, whereas vilt-b32-finetuned-vqa maintains better accuracy at even lower ranks. The latency per sample for blip-vqa-base at k = 0.8, k = 0.6, and k = 0.4 is 0.0442s, 0.0435s, and 0.0429s, respectively, while for vilt-b32-finetuned-vqa, it is 0.0148s, 0.0136s, and 0.0136s at the same rank reduction levels. We observe that low-rank approximation leads to lower latency per sample compared to quantization and pruning. We implement low-rank approximation by iterating over each linear layer, computing a reduced rank from a given ratio, and performing SVD on the weight matrix. Only the top singular values and vectors are kept to reconstruct an approximated weight, which replaces the original.
Figure 17: Evaluation of various VLMs across multiple multimodal benchmarks, including MMMU [205], MMB [111], ScienceQA [120], AI2D [80], SEED [91], and MathVista [119]. The average score represents the mean performance across all benchmarks. Data for this plot is sourced from [19].
Figure 18: Evaluation of VLMs of varying sizes (256M, 500M, 2.2B, and 80B) across multiple multimodal benchmarks, including OCRBench [112], TextVQA [156], DocVQA [132], ChartQA [131], AI2D [80], MathVista [119], MMStar [19], MMMU [205], and ScienceQA [120]. Data for this plot is sourced from https://huggingface.co/blog/smolervlm.
7.2 Evaluating VLM Performance on Benchmark Multimodal Datasets
We present the performance of various VLMs with differing parameter counts across widely used benchmarks to enable comparative analysis. Figure 17 illustrates the average scores of 14 different VLMs evaluated on 6 multimodal benchmarks. The average score is depicted in the figure. Interestingly, the results show that a higher parameter count does not necessarily translate to better accuracy. For example, InternLM2-XC2, with 7B parameters, achieves an impressive average score of 72, while Sphinx-X-MoE, despite having 57B parameters, achieves a lower average score of just 60.1. Similarly, comparatively smaller models like TinyLLaVA and Yi-VL, with 3B and 6B parameters respectively, achieve competitive average scores around 56. These findings suggest that the performance of a VLM depends more on the optimization method and design choices than on model size alone, which is an important insight for selecting or designing models suitable for deployment on edge devices with limited computational resources. We also report the accuracy (%) of very small VLMs such as SmolVLM (256M and 500M), medium-sized models like SmolVLM (2.2B), and large-scale models such as Idefics (80B) across nine multimodal benchmarks. As shown in Figure 18, SmolVLM with 2.2B parameters consistently achieves the highest accuracy across eight benchmarks, followed by the 500M and then the 256M models. Interestingly, Idefics with 80B parameters yields the lowest performance across most benchmarks. In terms of throughput, the 256M model delivers the highest, followed by the 500M, 2.2B, and finally the 80B model. A higher throughput indicates better efficiency. These results highlight a critical insight: achieving optimal performance on edge or pervasive IoT devices requires a careful balance between accuracy and efficiency, tailored to the specific constraints of the deployment hardware, such as memory and GPU availability.

### v3

@@ 3.1.2 Prompt Tuning
In prompt tuning [75], task-specific "soft prompts" are added to the input data. These are opaque learnable tensors concatenated with the input embedding as shown in figure 9. Throughout training, only the prompts are modified for the specified task. This strategy is extremely beneficial for multitasking and continual learning.
@@
[161] introduce Knowledge-guided Context Optimization (KgCoOp) to enhance the generalization capability of prompt tuning for VLMs. Their strategy reduces the discrepancy between learned soft prompts and hard prompts, thus reducing the likelihood of forgetting important general knowledge. By integrating this optimization constraint with contrastive learning, their method improves performance on unseen classes while also maintaining computational efficiency. [154] propose a dual-modality prompt tuning (DPT) approach that optimizes both visual and textual prompts simultaneously to adapt vision-language models to downstream tasks. They also introduce class -aware visual prompts, generated by cross-attention between textual features and image tokens, to improve task-specific feature extraction. To address the misalignment between text and image embeddings in VLMs, [17] develop DAPT to optimize feature distributions by maximizing inter-class separation and minimizing intra-class variability. This approach improves generalization in few-shot learning and domain adaptation.
3.1.3 Adapter-based Methods
@@
$X\in\mathbb{R}^{a\times a}\rightarrow W_{d}X\in\mathbb{R}^{a\times b}\rightarrow W_{u}(W_{d}X)\in\mathbb{R}^{a\times a}$ (12)
@@ 3.1.4 Prefix Tuning
Prefix Tuning [80] introduces task-specific continuous vectors into each layer of the transformer architecture. Prefixes are abstract representations concatenated with input embeddings in each layer. [130] introduce a mapping network that transforms visual features into token embeddings and employs prefix tuning to to refine query  representations in each transformer layer for open-ended medical visual question answering (VQA). [129] propose a vision-language storytelling framework that integrates a trainable transformer mapping network and a multimodal contrastive objective to enhance coherence in generated narratives. Prefix tuning is employed to align CLIP-extracted visual features with a frozen GPT-2 [119] model, facilitating efficient vision-language alignment. Prefixes are introduced in personalized image captioning to map visual features with user context, enabling a frozen GPT-2 model to provide personalized captions [148].
@@ 4 Runtime Optimizations
Runtime optimization refers to techniques that dynamically adjust computations during  inference to improve computational efficiency, memory use, and energy consumption without requiring retraining. Unlike static optimizations such as pruning or quantization, runtime methods adapt model execution on-the-fly based on input complexity, hardware constraints, or task-specific requirements. VLMs are computationally intensive, necessitating significant processing for multimodal interactions. Their deployment on  resource-constrained devices is challenging due to the high memory, power, and latency demands. Runtime optimization mitigates these issues by reducing unnecessary computations and accelerating inference.
@@ 4.2 Test-Time Adaptation (TTA)
$p(y\mid x)=\arg\max_{y}\mathbb{E}_{q(\theta)}[p(y\mid x;\theta)]$ (13)
@@ 5 Privacy-preserving Distributed VLM
Privacy-preserving distributed VLM ensures that model training occurs across multiple devices or servers without sharing raw data. The key benefits include enhanced data security and a decreased risk of data leakage. Below, we discuss key techniques such as Federated Learning (FL) 5.1 and Split Learning 5.2.
@@ 6.5 MiniVLM
6.6 LightCLIP
6.7 NanoVLMs
6.8 Moondream
6.9 Xmodel-VLM
Furthermore, Table 2 outlines various frameworks and libraries and highlights key features of each.
@@ 6.9 Xmodel-VLM
DeepSpeed  [^1] https://www.deepspeed.ai/ Zero and sparse training optimizations Distributed fine-tuning and model parallelism Large VLMs (e.g., LLaVA variants, Flamingo, MiniGPT-4, etc.) Not directly
@@ 7 Analysis and Insights
In this section, we analyze the impact of efficiency-driven techniques on accuracy, latency, and memory consumption. Our evaluation is based on two models: 1 blip-vqa-base [77] ($\sim 385M$ parameters) and 2 vilt-b32-finetuned-vqa [71] ($\sim 87M$ parameters). All results are reported using an NVIDIA RTX A6000 GPU. Notably, the observed trends remain consistent across all models, provided the parameter count falls within the range of these two models. This section provides insights specifically for pre-deployment techniques such as quantization, pruning, and low-rank approximation. We focus on these methods because they are fundamental for optimizing models before deployment, directly impacting model size and performance. Although efficient fine-tuning, runtime optimizations, and privacy-preserving methods are important, they are often model- or task-specific and involve different challenges and considerations that extend beyond the scope of this section. Our analysis provides insights into the accuracy vs. efficiency trade-off, emphasizing the importance of selecting the right technique based on the specific requirements of the downstream application.
@@
Quantization (figure 14) has a notable impact on accuracy, latency, and memory consumption on both blip-vqa-base and vilt-b32-finetned-vqa. For blip-vqa-base, transitioning from FP32 to INT8 results in a 0.85% accuracy drop for 50 samples whereas vilt-b32-finetned-vqa experiences an 8.07% reduction indicating that ViLT is more sensitive to INT8 quantization. We also increase the number of samples to 125 to observe the effect on accuracy and memory consumption. Increasing the sample size from 50 to 125 causes a 1.36% accuracy drop in FP32 and a 1.54% drop in INT8 for blip-vqa-base, while vilt-b32-finetned-vqa sees a 7.1% decrease in FP32 and a 0.09% reduction in INT8. However, the accuracy drop for vilt-b32-finetuned-vqa from FP32 to INT8 is steeper (4.6% on average) compared to blip-vqa-base (0.94%), suggesting that vilt-b32-finetuned-vqa benefits more from higher precision, but suffers greater degradation from aggressive quantization. In terms of memory consumption, quantization significantly reduces memory requirements across both models. For blip-vqa-base, moving from FP32 to INT8 results in a 53.8% reduction (747.8MB to 345.2MB) for 50 samples and a 43.1% reduction (800MB to 455MB) for 125 samples, while vilt-b32-finetuned-vqa achieves a 57.2% reduction (586.3MB to 251.1MB) for 50 samples and 44.2% (664.4MB to 370.6MB) for 125 samples, demonstrating that INT8 provides substantial memory savings, with vilt-b32-finetuned-vqa benefiting slightly more. Across all quantization levels, blip-vqa-base consistently consumes more memory than vilt-b32-finetuned-vqa, with a difference of 161.5MB in FP32 (50 samples) and 94.1MB in INT8 (50 samples). The blip-vqa-base maintains a more stable accuracy across sample sizes but consistently requires more memory. The latency per sample remains 0.0877s, 0.0757s, and 0.0701s across FP32, FP16, and INT8 for blip-vqa-base and 0.0330s, 0.0321s, and 0.0311s for vilt-b32-finetuned-vqa, with INT8 being the fastest in both cases. For FP32, we run the models with full 32-bit precision. For FP16, mixed precision is enabled using autocasting to boost speed while maintaining accuracy. For INT8, a copy is dynamically quantized on its linear layers to shrink the model and speed up inference.
@@
Pruning (figure 15) significantly impacts accuracy and memory consumption across both models. For blip-vqa-base, accuracy drops by 0% at 20% pruning, 4.24% at 40% pruning and 11.9% at 60% pruning for 50 samples. The vilt-b32-finetuned-vqa model follows a similar trend, but at 60% pruning, accuracy drops to 0%, indicating extreme sensitivity to high sparsity levels. In terms of memory savings, blip-vqa-base reduces by 16.3% at 20% pruning, 22.3% at 40% pruning, and 27% at 60% pruning (50 samples), while vilt-b32-finetuned-vqa achieves reductions of 1.7%, 15.4%, and 18.5%, respectively. For 125 samples, blip-vqa-base saves 12.3%, 20.7%, and 24.9%, whereas vilt-b32-finetuned-vqa reduces by 4.2%, 13.4%, and 21%. The results show that vilt-b32-finetuned-vqa maintains better accuracy at lower pruning levels but is highly sensitive to aggressive pruning. The latency per sample for blip-vqa-base at 20%, 40%, and 60% pruning is 0.0569s, 0.0561s, and 0.0559s, respectively, while for vilt-b32-finetuned-vqa, it is 0.0148s, 0.0141s, and 0.0139s at the same pruning levels. We observe that pruning results in faster inference per sample compared to quantization. We implement pruning by scanning all model modules to find linear layers and then apply global unstructured L1 pruning with a set sparsity ratio. After pruning, the extra reparameterization is removed, making the pruned model ready for inference.
@@
Low-rank approximation (figure 16) significantly affects accuracy and memory consumption across both models. Decreasing the rank from k = 0.8 to k = 0.4 increases compression, leading to lower memory consumption and latency. However, this also results in a drop in accuracy. For blip-vqa-base, accuracy decreases by 15.8% at k = 0.8, 27.8% at k = 0.6, and drops to 0% at k = 0.4 for 50 samples. The vilt-b32-finetuned-vqa exhibits a more gradual decline, with accuracy dropping by 4% at k = 0.8, 6.2% at k = 0.6, and 27.1% at k = 0.4, indicating that vilt-b32-finetuned-vqa is more resilient to LRA compared to blip-vqa-base. In terms of memory savings, blip-vqa-base reduces memory consumption by 140 MB at k = 0.8, 160.7 MB at k = 0.6, and 208.6 MB at k = 0.4 (50 samples), while vilt-b32-finetuned-vqa achieves reductions of 7.2 MB, 63.2 MB, and 86.4 MB, respectively. For 125 samples, blip-vqa-base saves 65.4 MB, 100.2 MB, and 199.3 MB, whereas vilt-b32-finetuned-vqa reduces by 64 MB, 83.8 MB, and 150 MB. This indicates that while low-rank approximation provides substantial memory savings, blip-vqa-base is highly sensitive to aggressive rank reduction, whereas vilt-b32-finetuned-vqa maintains better accuracy at even lower ranks. The latency per sample for blip-vqa-base at k = 0.8, k = 0.6, and k = 0.4 is 0.0442s, 0.0435s, and 0.0429s, respectively, while for vilt-b32-finetuned-vqa, it is 0.0148s, 0.0136s, and 0.0136s at the same rank reduction levels. We observe that low-rank approximation leads to lower latency per sample compared to quantization and pruning. We implement low-rank approximation by iterating over each linear layer, computing a reduced rank from a given ratio, and performing SVD on the weight matrix. Only the top singular values and vectors are kept to reconstruct an approximated weight, which replaces the original.